OpenStudy (anonymous):

Which equation is not true for all values of 0? a. csc 0 = sec 0 / tan 0 cos 0 b. sin ^2 0+cos^2 0 = sec^2 0 (1-sin^2 0) c. cot 0 - tan 0 = cos 0 csc 0 - sin 0 sec 0 d. sin ^2 0 + cos ^2 0 + tan ^2 0 = sec^2 0

OpenStudy (anonymous):

@aum :)

OpenStudy (aum):

Is it zero or the Greek letter phi \(\phi\) or \(\theta\)?

OpenStudy (anonymous):

The one with a line through it. Sorry, I forgot to add that.

OpenStudy (aum):

It is phi.

OpenStudy (aum):

Is the first choice \[ \csc(\phi) = \frac{\sec(\phi)}{\tan(\phi)\cos(\phi)}~~\text{or } \\ \csc(\phi) = \frac{\sec(\phi)}{\tan(\phi)}\cos(\phi) \]

OpenStudy (anonymous):

follow me savanah